OpenStudy (anonymous):

equation that I need help with is 8 over 3 plus sqrt 7

OpenStudy (dumbcow):

\[\frac{8}{3+\sqrt{7}}\]

OpenStudy (dumbcow):

is that it

OpenStudy (dumbcow):

well if it is multiply top and bottom by conjugate \[\frac{8}{3+\sqrt{7}}*\frac{3-\sqrt{7}}{3-\sqrt{7}}\] \[=\frac{24-8\sqrt{7}}{2}\] factor out a 2 on top which cancels the 2 on bottom leaving \[= 12-4\sqrt{7}\]

OpenStudy (anonymous):

wonderful! how did you come up with the 8, though? is it that you must include the same number you multiply with in the top equation?

OpenStudy (anonymous):

and how in the heck did you get 2?

OpenStudy (anonymous):

oh, never mind. thank you!

OpenStudy (dumbcow):

by distributing on top, FOIL on bottom your welcome :)

OpenStudy (anonymous):

quick question- for these types of problems, will the denominator always be 2?